I am having a lot of difficulty with this problem.
I hope the modified square root symbol is clear. The square root symbol is covering the 2x + 7.
Solve the equation:
_____
√2x + 7 - (x + 2) = 0

Answer by stanbon(75887)

√(2x + 7) - (x + 2) = 0
----
sqrt(2x+7) = (x+2)
Square both sides to get:
2x+7 = x^2+4x+4
---
Modify:
x^2 + 2x - 3 = 0
---
Factor:
(x+3)(x-1) = 0
---
Possible answers: x = -3 or x = 1
---
Checking the possible answers in the original equation:
sqrt(2x+7) = (x+2)
x= -3 ???
sqrt(2(-3)+7) =(-3+2)
sqrt(1) = -1
No
----------------
x = 1 ???
sqrt(2x+7) = (x+2)
sqrt(2(1)+7) = (1+2)
sqrt(9) = 3
3 = 3
Yes
=======================
Cheers,
Stan H.
===============